Return  Young

Direct(Shift JIS)

、。，．・：；？！゛゜´｀¨＾￣＿ヽヾゝゞ〃仝々〆〇ー―‐ ／＼〜‖｜…‥‘’“”（）〔〕［］｛｝〈〉《》「」『』【】＋−±×÷＝≠＜＞≦≧∞∴♂♀°′″℃￥＄¢£％＃＆＊＠§ ☆★○●◎◇◆□■△▲▽▼※〒→←↑↓〓ぁあぃいぅうぇえぉおかがきぎくぐけげこごさざしじすずせぜそぞただちぢっつづてでとどなにぬねのはばぱひびぴふぶぷへべぺほぼぽまみむめもゃやゅゆょよらりるれろゎわゐゑをんァアィイゥウェエォオカガキギクグケゲコゴサザシジスズセゼソゾタダチヂッツヅテデトドナニヌネノハバパヒビピフブプヘベペホボポマミムメモャヤュユョヨラリルレロヮワヰヱヲンヴヵヶ０１２３４５６７８９ＡＢＣＤＥＦＧＨＩＪＫＬＭＮＯＰＱＲＳＴＵＶＷＸＹＺａｂｃｄｅｆｇｈｉｊｋｌｍｎｏｐｑｒｓｔｕｖｗｘｙｚ ΑΒΓΔΕΖΗΘΙΚΛΜΝΞΟΠΡΣΤΥΦΧΨΩ αβγδεζηθικλμνξοπρστυφχψω АБВГДЕЁЖЗИЙКЛМНОПРСТУФХЦЧШЩ ЪЫЬЭЮЯ абвгдеёжзийклмнопрстуфхцчшщ ъыьэюя ∈∋⊆⊇⊂⊃∪∩∧∨¬⇒⇔∀∃∠⊥⌒∂∇≡≒≪≫√∽∝∵∫∬ Å‰♯♭♪†‡¶◯

HTML

% ‰ ¡ @ & ¢£ ¥ § † ‡ ¬ \begin{matrix}c \\ ◯ \end{matrix} ◯^{R} ^{TM} ^{ ¨ a} ^{ ¯ ∘} ± ´ μ ¶ · ¸ ≪ ≫ ¿

0 ^{1} ^{2} ^{3} 1/41/23/41/83/85/87/81/32/31/52/53/54/5

À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì Í Î Ï Ñ Ò Ó Ô Õ Ö Ø Ù Ú Û Ü Ý à á â ã ä å æ ç è é ê ë ì í î ï ñ ò ó ô õ ö ø ù ú û ü ý ß ÿ

• ⋯ ′ ′′ ← ↑ → ↓ ↔ ↕ ↖ ↗ ↘ ↙ ⇒ ⇔ A B Γ Δ E Z H Θ I K Λ M N Ξ O Π P Σ T Υ Φ X Ψ Ω α β γ δ ϵ ζ η θ ι κ λ μ ν ξ o π ρ ς σ τ υ φ χ ψ ω ϑ ϖ ∀ ∂ ∃ ∅ ∇ ∈ ∉ ∋ ∏ ∑ − ∗ √ ∝ ∞ ∠ ∧ ∨ ∩ ∪ ∫ ∫∫ ∮ ∴ ∵ ∼ ≅ ≈ ≠ ≡ ≤ ≥ ⊂ ⊃ ⊄ ⊆ ⊇ × ÷ ⊕ ⊗ ⊥ · ℧ < > ′′ Œ œ S̆ s̆ Ÿ ˆ ˜ – — ‘ ’ , “ ” ,, ⟨ ⟩ ▲ ▶ ▼ ◀ ■ □ ○●♀♂♠ ♣ ♡ ♦ ⋆ ☆

Top Packages for Young Tableau

1. young.sty (by Jörg Knappen)

\usepackage{young}
\begin{Young}
..  & ..  & ..  \cr
..  & ..  & ..  \cr
\end{Young}
\begin{Young}
 $a_1$ & $a_2$ & $a_3$ & $a_4$ \cr
 $b_1$ & $b_2$ & $b_3$ \cr
 $c_1$ & $c_2$ & $c_3$\cr
 $d_1$ & $d_2$\cr
\end{Young}
\begin{Young}
  p  &  q &  r & s \cr
  pa & pb & pc \cr
  qa & qb & qc \cr
  ra & rb \cr
\end{Young}

\begin{matrix}a_{1} & a_{2} & a_{3} & a_{4} \\ b_{1} & b_{2} & b_{3} \\ c_{1} & c_{2} & c_{3} \\ d_{1} & d_{2} \end{matrix}

p q r s pa pb pc qa qb qc ra rb

2. youngtab.sty (by Volker Böchers and Stefan Gieseke)

\usepackage[..]{youngtab}
vcentermath        \Yvcentermath=0 or 1
noautoscale        \Yautoscale=0 or 1
stdtext
enableskew

\Yboxdim= (valid only when autoscale is off) \Ylinethick= (default: 0.3pt) \Yinterspace= (default: 0ex plus .3ex)

\def\I{ii}
$\yng(1,1)\oplus\yng(1) = \yng(2,1) + \cdots$
{\tiny$\yng(1,1)\oplus\yng(1) = \yng(2,1) + \cdots$}
\young(\I bcd,efg,hij,k) \hfill \young(abcd,:bc,::d)

⊕ =  + ⋯  ⊕ =  + ⋯  ii b c d ; e f g ; h i j ; k a b c d ; b c ; d